# SpiralShard: Highly Concurrent and Secure Blockchain Sharding via Linked Cross-shard Endorsement

### v1

@@ VI-B Parameter Settings
The parameter settings are shown in Table I. By default, we choose the shard size of around 100 for SpiralShard, which is practical and used by some previous works [46, 23, 29, 21, 17]. According to Table I, SpiralShard significantly reduces the shard size under the negligible failure probability, increasing the number of shards under the same network size. As shown in Table II, we provide a reference to update group sizes with increased network sizes when developers have set different shard sizes. For example, given a shard size of 100 and network sizes, the system follows the second column to set endorsement group sizes. Additionally, when the network scale is fixed, we recommend that developers of latency-sensitive systems increase the shard size to reduce the size of endorsement groups, thereby achieving the goal of trading throughput for lower latency (see Section VI-E).
TABLE I: Parameter settings in Harmony and SpiralShard.
Network Size 1260 2000 3000 4200
@@
Shard Size 420 500 600 600
# of Shards 3 4 5 7
Failure Probability ($\cdot 10^{-6}$) 3.9 4.3 0.8 3.1
@@
Shard Size 105 100 100 100
# of Shards 12 20 30 42
Size of Endorsement Group 4 5 6 6
# of Endorsement Group 3 4 5 7
Failure Probability($\cdot 10^{-6}$) 6.5 4.9 2.1 5.3
TABLE II: Reference to endorsement group sizes of SpiralShard.
Network Size Shard Size 100 120 150 200 300
500 $\sim$ 699 3 3 2 2 1
700 $\sim$ 899 4 3 3 2 2
900 $\sim$ 1499 4 4 3 2 2
1500 $\sim$ 2099 5 4 3 3 2
2100 $\sim$ 2299 5 4 4 3 2
2300 $\sim$ 5299 6 5 4 3 2
5300 $\sim$ 5899 7 5 4 3 2
5900 $\sim$ 8299 7 6 5 4 3
8300 $\sim$ 10000 8 6 5 4 3
VI-C Throughput
We compare the average throughput in terms of transactions per second (TPS) of SpiralShard and the baseline at different network sizes. As shown in Figure 6a, SpiralShard achieves more than 9.5$\times$ throughput. Moreover, SpiralShard achieves up to 19.5$\times$ throughput gain at a network size of 3,000. The main reasons behind this are three-fold. First, SpiralShard achieves reduced shard size, resulting in a faster intra-shard consensus. Second, the larger number of shards in SpiralShard brings higher transaction concurrency. Third, the pipelining mechanism allows each shard to continuously produce blocks without being halted by the slower shards.
VI-D Latency
We compare the average transaction confirmation latency of SpiralShard and the baseline under different network sizes. Transaction confirmation latency here is the duration from when a transaction starts to be processed until the transaction is confirmed, the same as Monoxide [45]. As shown in Figure 6b, SpiralShard has longer latency (up to 1.9$\times$ at a network size of 4,200). By comparing Figure 6a and Figure 6b, we find that although the throughput of SpiralShard is greatly improved, the transaction confirmation latency is not too large. The main reason is that, although SpiralShard requires each block to be sequentially endorsed by other shards, the significantly reduced shard size speeds up intra-shard consensus, shortening the latency for each step of the LCE, resulting in an acceptable overall latency.
@@ VI-D Latency
VI-E Performance under Various Shard Sizes
@@
(a) Total TPS
(b) Confirmation latency
Figure 8: Performance results under fixed endorsement group size.
VI-F Performance under Same Endorsement Group Size
In this experiment, we set the endorsement group size as 2 and evaluated the performance at different network sizes. The experimental configurations are shown in Table IV. As illustrated in Figure 8b, SpiralShard can achieve a transaction confirmation latency similar to the baseline protocol. The intuition is that it takes less time to finalize a block due to the smaller endorsement group size. More importantly, as shown in Figure 8a, SpiralShard still achieves more than 3.8$\times$ throughput improvement. The reason is that SpiralShard can be configured with smaller and more shards.
TABLE IV: Parameter settings in SpiralShard under fixed endorsement group size.
Network Size 1260 2000 3000 4200
Shard Size 210 250 300 300
# of Shards 6 8 10 14
# of Endorsement Group 3 4 5 7
VI-G Performance with More Malicious Nodes
We evaluate the performance of SpiralShard in the presence of $1/3$ fraction of total malicious nodes (with $15\%$ fraction of Byzantine nodes) under the 800 network size. The results are shown in Table V. Specifically, with such a high fraction of total malicious nodes, the baseline can only configure one shard to ensure the system’s security, while SpiralShard can configure 8 shards. More importantly, SpiralShard achieves 49x higher throughput than the baseline protocol. However, the latency of SpiralShard is higher. The main reason is that SpiralShard adopts a large endorsement group, extending the latency for a block to be finalized.
TABLE V: Configuration and evaluation results under network size of 800.
Baseline SpiralShard
# of Shards 1 8
Shard Size 800 100
# of Endorsement Group - 1
Failure Probability($\cdot 10^{-6}$) 0 0.04
Total TPS 34 1660
Confirmation Latency (s) 119.54 157.9
VI-H Performance with a Slow Shard
In this experiment, we evaluate the performance of SpiralShard in the presence of one slow shard under the configuration shown in Table V. Specifically, we partition the network inside a shard for a while to simulate the situation where a shard cannot perform consensus due to network problems during a certain period, similar to previous work [43].
As illustrated in Figure 9, we start the network partition in a shard at 75s and continue for 60s. It is shown that the number of finalized blocks gradually decreases during this period. The reason is that since our LCE protocol requires sequential endorsement between shards within the endorsement group, a slow shard will halt the LCE’s processing within the endorsement group. This results in a decreasing number of blocks that can be finalized. Additionally, to fully reflect the impact of slow shard within the endorsement group, we have configured only one endorsement group. After the network partition period, the consensus process for that slow shard resumes, and the number of finalized blocks increases rapidly. This is because, in the LCE protocol, when the newest block collects enough endorsements, all of it and its parent blocks are finalized. Therefore, those blocks not finalized during the network partition period will be finalized together rapidly after the consensus is restored.
VI-I Performance During an Epoch Transition
@@
In this section, we evaluate the performance during an epoch transition. Specifically, as the system consistently finalizes blocks, we trigger an epoch transition at $T=5$. As shown in Figure 10a, there is a sharp drop in the throughput following the onset of the epoch transition. This is primarily due to state synchronization conducted by nodes before joining new shards. The subsequent peak in throughput stems from the new epoch’s finalizing prepared blocks from the previous epoch. As observed in Figure 10b, there is a peak in latency after the conclusion of the epoch transition. The main reason is that blocks generated toward the end of an epoch necessitate sequential endorsement from scratch in the new epoch. These findings underscore that system performance experiences an impact during epoch transitions. However, since epoch transitions occur only once daily, aligning with existing works [46, 29, 21, 31], the performance remains unaffected most of the time.

### v2

@@ VII-B Parameter Settings
By default, we choose the shard size of around 100 for SpiralShard, which is practical and used by some previous works [rapidchain, omniledger, pyramid, sgxSharding]. Indeed, SpiralShard can be configured with even more amount of smaller shards while maintaining the same level of security guarantees. According to Table II, SpiralShard significantly reduces the shard size under the negligible failure probability, increasing the number of shards under the same network size. It is worth noting that GearBox fails to successfully configure a large number of secure shards. This is primarily because, after the shard-scaling process, many shards still contain a fraction of malicious nodes exceeding their fault tolerance threshold, resulting in widespread shard configuration failures.
TABLE II: Parameter settings in Harmony and SpiralShard.
Network Size 800 1260 2000 3000 4200
@@
Shard Size 400 420 500 600 600
# of Shards 2 3 4 5 7
Failure Probability ($\cdot 10^{-6}$) 0.9 3.9 4.3 0.8 3.1
GearBox
Shard Size 69 76 77 82 82
# of Shards 5 8 11 14 14
@@
Shard Size 100 105 100 100 100
# of Shards 8 12 20 30 42
Size of Endorsement Group 4 4 5 6 6
# of Endorsement Group 2 3 4 5 7
Failure Probability($\cdot 10^{-6}$) 0.1 6.5 4.9 2.1 5.3
VII-C Throughput
[…]
Notably, SpiralShard achieves considerable performance improvements even with a higher cross-shard transaction ratio compared to the baseline. This indicates that the performance advantages of SpiralShard’s unique system design outweigh the negative impacts of increased cross-shard transaction ratios. It also clearly suggests that, if the cross-shard transaction ratio between SpiralShard and the baseline are artificially controlled to be equal (possibly by employing some mechanisms to reduce the cross-shard transaction ratio [LBCHAIN, brokerchain]), SpiralShard could achieve even better performance.
VII-D Latency
[…]
@@ VII-D Latency
VII-E Performance under Various Shard Sizes
@@
Based on the above experimental results, SpiralShard provides developers with flexible sharding configurations to suit different application scenarios. For instance, some scenarios involve large volumes of transaction data but have low demands for immediate processing. Typical examples include cross-border payments and settlements, which can tolerate delays ranging from a few minutes to hours but require high throughput to handle large-scale transaction demands, or use cases such as supply chain tracking and IoT device logging. For these scenarios, developers can configure a larger number of smaller shards (e.g., the left area in Figure 7a and 7b) to achieve higher throughput. On the other hand, for scenarios that impact user experience, such as blockchain-based login verification, access control, or on-chain voting and governance—applications that are relatively sensitive to latency—developers can configure larger shards (e.g., the right area in Figure 7a and 7b), trading some throughput for reduced latency. More importantly, given the flexibility of SpiralShard’s sharding configurations, developers could even design mechanisms to dynamically adjust the sharding configuration within the system, thereby adapting to changing application demands in real time. We consider this direction as a potential avenue for future research.
VII-F Overhead of LCE Protocol
Figure 8: Consensus time comparsion.
We now evaluate the impact of LCE protocol’s communication overhead on consensus performance. Consensus time (or speed) serves as a fundamental metric for assessing system performance, as it directly influences both throughput and transaction latency. Specifically, we isolate the cost introduced by the LCE’s broadcasting and measure the time consumed by intra-shard consensus under normal conditions. As illustrated in Figure 8, the communication overhead caused by cross-shard endorsement is negligible ($\approx 1\%$). This is because, propagating cross-shard transactions already requires each shard to broadcast block headers to other shards. In SpiralShard, cross-shard endorsement simply reuses these block headers, adding only a few integer-sized data fields (see Figure 3).
VII-G Performance During an Epoch Transition
@@
In this section, we evaluate the performance during an epoch transition (under 1,260 nodes). Specifically, as the system consistently finalizes blocks, we trigger an epoch transition at $T=5$. As shown in Figure 9a, there is a sharp drop in the throughput following the onset of the epoch transition. This is primarily due to state synchronization conducted by nodes before joining new shards. The subsequent peak in throughput stems from the new epoch’s finalizing prepared blocks from the previous epoch. As observed in Figure 9b, there is a peak in latency after the conclusion of the epoch transition. The main reason is that blocks generated toward the end of an epoch necessitate sequential endorsement from scratch in the new epoch. These findings underscore that system performance experiences an impact during epoch transitions. However, since epoch transitions occur only once daily, aligning with existing works [rapidchain, omniledger, pyramid, li2022jenga], the performance remains unaffected most of the time.
@@
VII-H Performance with Slow Shards
In this experiment, we evaluate the performance of SpiralShard in the presence of slow shards. Specifically, we partition the network inside a shard for a while to reflect the situation such as temporary consensus failures caused by potential problems like silent attacks launched by malicious nodes.
As illustrated in Figure 10, we start the network partition in a shard at 75s and continue for 60s. It is shown that the number of finalized blocks gradually decreases during this period. The reason is that since our LCE protocol requires sequential endorsement between shards within the endorsement group, a slow shard will halt the LCE’s processing within the endorsement group. This results in a decreasing number of blocks that can be finalized. Additionally, to fully reflect the impact of slow shard within the endorsement group, we have configured only one endorsement group. After the network partition period, the consensus process for that slow shard resumes, and the number of finalized blocks increases rapidly. This is because, in the LCE protocol, when the newest block collects enough endorsements, all of it and its parent blocks are finalized. Therefore, those blocks not finalized during the network partition period will be finalized together rapidly after the consensus is restored. Noted that SpiralShard’s safety and liveness remain intact. In the silent attack case, no conflicting blocks are ever finalized, preserving safety. In addition, once the affected shard recovers, all pending blocks are finalized in sequence, showing that liveness is only temporarily hindered but not permanently lost.
VII-I Performance under Network Fluctuation
In this section, we evaluate the performance under adversarial network fluctuation to assess the robustness (under 800 nodes with 1 endorsement group). This evaluation can reflect the situation such as attacters delaying the network. Specifically, we use different probability distributions to model the additional delay to the propagation of each message. These distributions are parameterized as follows: a uniform distribution with a range of 100 to 300 ms, a normal distribution with a mean ($\mu$) of 200 ms and a standard deviation ($\sigma$) of 50 ms, and a log-normal distribution where the logarithm of the delay follows a normal distribution with $\mu=5.268$ and $\sigma=0.246$. In contrast, the default case serves as a baseline with no additional propagation delay applied.
As shown in Figure 11a and Figure 11b, the additional delay results in a reduction in system throughput and an increase in system latency compared to the default network environment. This is primarily because the extra delay extends the time required to reach consensus. Nevertheless, the impact on system performance remains within acceptable bounds, with throughput loss and latency increase constrained to within 10% and 12%, respectively. These results also show that SpiralShard’s safety and liveness are not affected. Under network delay attacks, the protocol continues to finalize blocks (albeit at a slightly reduced rate), ensuring the system does not stall and all valid transactions eventually confirm
(a) Total TPS
(b) Confirmation latency
Figure 11: Performance with different message propagation delay.
VIII Conclusion
[…]
[Uncaptioned image] You Lin is currently a master candidate with Department of Computer Science and Engineering, Southern University of Science and Technology. He received his B.E. degree in computer science and technology from Southern University of Science and Technology in 2021. His research interests are mainly in blockchain, network economics, and consensus protocols.
[Uncaptioned image] Mingzhe Li is currently a Scientist with the Institute of High Performance Computing (IHPC), A*STAR, Singapore. He received his Ph.D. degree from the Department of Computer Science and Engineering, Hong Kong University of Science and Technology in 2022. Prior to that, he received his B.E. degree from Southern University of Science and Technology. His research interests are mainly in blockchain sharding, consensus protocol, blockchain application, network economics, and crowdsourcing.
[Uncaptioned image] Jin Zhang is currently an associate professor with Department of Computer Science and Engineering, Southern University of Science and Technology. She received her B.E. and M.E. degrees in electronic engineering from Tsinghua University in 2004 and 2006, respectively, and received her Ph.D. degree in computer science from Hong Kong University of Science and Technology in 2009. Her research interests are mainly in mobile healthcare and wearable computing, wireless communication and networks, network economics, cognitive radio networks and dynamic spectrum management.